OpenStudy (anonymous):

convert 8.987 decades to min

ganeshie8 (ganeshie8):

1 decade = 10 years

ganeshie8 (ganeshie8):

1 year = how many days ?

OpenStudy (anonymous):

365

ganeshie8 (ganeshie8):

good, no ambiguity there, :)

OpenStudy (gitface):

Unless you want to take into account leap years, which gets messy.

ganeshie8 (ganeshie8):

dimensional analysis now

OpenStudy (anonymous):

yeah @ganeshie8 so whats after 1 year=365 days

ganeshie8 (ganeshie8):

\(8.987 \) decades

OpenStudy (anonymous):

nvm i got it now i'll let you know my final answer when i figure it out

ganeshie8 (ganeshie8):

\(8.987 \) \( decades\) x \(\frac{10 years}{1 decade}\)

ganeshie8 (ganeshie8):

\(8.987 \) \( decades\) x \(\frac{10 years}{1 decade}\) x \(\frac{365 days}{1 year}\)

ganeshie8 (ganeshie8):

\(8.987 \) \( decades\) x \(\frac{10 years}{1 decade}\) x \(\frac{365 days}{1 year}\) x \(\frac{24hours}{1 day}\)

OpenStudy (anonymous):

yeah sooo my final answer is 47235672

ganeshie8 (ganeshie8):

\(8.987 \) \( decades\) x \(\frac{10 years}{1 decade}\) x \(\frac{365 days}{1 year}\) x \(\frac{24hours}{1 day}\) x \(\frac{60 mnts}{1 hour}\)

ganeshie8 (ganeshie8):

all units cancel out, except mnts ha ?, so lets see if i get the same answer as urs :)

OpenStudy (anonymous):

okay and yeah i did cancle all the units except min

ganeshie8 (ganeshie8):

\(8.987 \) \( \cancel{decades}\) x \(\frac{10 \cancel{years}}{1 \cancel{decade}}\) x \(\frac{365 \cancel{days}}{1 \cancel{year}}\) x \(\frac{24\cancel{hours}}{1 \cancel{day}}\) x \(\frac{60 mnts}{1 \cancel{hour}}\)

ganeshie8 (ganeshie8):

we just multiply together the numbers nw

ganeshie8 (ganeshie8):

8.987 x 10 x 365 x 24 x 60

OpenStudy (anonymous):

all this work you showing is very confusing i know how to do it but is my answer right?

ganeshie8 (ganeshie8):

i get 47235672

ganeshie8 (ganeshie8):

good work !

ganeshie8 (ganeshie8):

one thing,

OpenStudy (anonymous):

good :) how can we make that number short

ganeshie8 (ganeshie8):

ya thats what ima say, we can convert it to scientific notation

OpenStudy (anonymous):

which will; be ?

ganeshie8 (ganeshie8):

\(47235672 minutes = 4.7235672 \times 10^7 minutes\)

ganeshie8 (ganeshie8):

4.7 is less than 47235672, in scientific notation, we always convert a number to a short nmber between (1-10) and use 10 powers for the exponent

ganeshie8 (ganeshie8):

i see u already knw these things and comfortable wid doing these ur way :)

OpenStudy (anonymous):

yeah but still what will be my final answer in scientific notation??

ganeshie8 (ganeshie8):

above i wrote it

ganeshie8 (ganeshie8):

after ur , which will; be ? response

OpenStudy (anonymous):

soo 4.7235672x10^7

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

thnaks :))

ganeshie8 (ganeshie8):

dont forget to put the unit minutes in the end :)

OpenStudy (anonymous):

yeah i know thanks